# What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality

### v2

@@ 2.3.1 From Signals to Perception
We first examine whether sensor-derived behavioral signals correspond to how users interpret collaboration. We call this a relationship. We call this relationship reflective interaction mapping. While subjective reports are post-hoc and summarized, we analyze objective indicators over task-level windows that match the granularity of users’ reflections. We do not compare specific events to single survey items but assess whether aggregate behavioral patterns (gaze overlap, turn-taking) are reflected in users’ overall experience. This avoids false precision and treats subjective reports as interpretive anchors rather than timestamped outcomes. For example, we ask whether shared gaze events align with a user’s sense of shared attention. Do balanced speech patterns correspond to reported conversational support? This connection motivates our first set of hypotheses, which examine whether subjective perceptions of collaboration are reflected in the observable sensor-based measures.
2.3.2 Interaction Traces and Performance
@@
2.3.3 Subjective Perception to Task Performance
Finally, we investigate whether users’ internal experiences correspond to observable group-level performance metrics. We refer to this relationship as a subjective influence on performance. Though not sensor-measured, states like presence and cognitive load shape how users engage with others and respond to collaborative demands. For example, lower cognitive load may support better coordination [15], leading to more balanced participation and faster task completion. This mapping tests hypotheses that investigate whether subjective experiences (own experience or the group as a whole) align with observable group outcomes.
@@ 2.4 Hypotheses
Together, the three relationships defined by S2S framework guide the structure of our hypotheses. We will focus our experimental design on testing the following hypotheses:
@@
H3.1: Higher individual perceived presence and perception of contribution are reflected in more frequent shared attention events among group members.
@@ 3.1 Participants
We recruited participants; data from 12 participants were discarded due to technical issues, resulting in participants being included in the analysis. Participants were allowed to form their own groups, or, if preferred, the research team randomly assigned them to a group. In total, 12 groups completed the study. In similar research on collaborative behavior involving small groups, the group is typically defined as having three or more members [64]; in our study, we formed a group with 4 members. Prior studies have revealed patterns of conversation, interaction, and coordination within groups of four in both desktop and virtual reality environments [70]. By adopting a group size of four participants, we aimed to create a richer collaborative environment, capturing the complexities of group behavior not as evident in smaller groups [3, 40]. Participants’ ages ranged from 21 to 42, with a mean age of 24. A summary of the participant demographics can be seen in Table 1. All participants provided verbal informed consent. Each participant had normal or corrected-to-normal vision. The institution’s ethics committee approved the study.
@@ 3.2 Materials
To capture sensor data on user interactions, we used the Meta Quest Pro [37]. The integrated sensors enabled audio recording, eye-tracking, six-degrees-of-freedom (6 DoF) simultaneous localization and mapping (SLAM) tracking, and MR capabilities. The collaborative app is built in Unity using Meta XR APIs to log audio, gaze, location, and object interactions [36]. For precise manipulation of virtual objects, we use the controller integrated with the Meta Quest Pro [37]. The interaction recording and tracking are validated by others [46]. Participants were invited to a shared lab room with a designated space cleared of materials to minimize distractions. They were informed that they could move freely within this area during the task.
@@ 3.3 Experimental Task
This section summarizes the cues, interaction, and feedback of our collaborative image-sorting task. The study was deliberately designed around a single collaborative task under one shared condition, without varying levels of network stressors or task types. Our goal was not to compare multiple experimental conditions but to examine the richness of group behavior in a naturalistic, unconstrained collaborative setting. Prior work has shown that tightly controlled comparisons can obscure the variability and fluidity of real-time collaboration in MR environments [70]. Instead, we focused on collecting high-resolution behavioral data and subjective reflections during a consistent, shared experience across groups. Each group of four participants completed one image-sorting task using the same images and categories. Participants were instructed to work together to reach a consensus on the grouping of each image.
@@ 3.3.1 Primary Task
Participants sorted 28 images from the OASIS dataset [31], which includes 900 validated images rated by 822 participants for pleasantness and arousal. Images were selected to span diverse emotions while avoiding distressing content. Participants sorted the selected images into one of six emotion categories randomly chosen from Russell’s circumplex model of affect [49]: angry, bored, relaxed, tense, pleased, or frustrated. Image sorting tasks have been shown to foster decision-making, communication, and social coordination by building shared mental models and group alignment [10]. We focus on this collaborative task marked by asynchronous, flexible participation, where participants can contribute and modify inputs independently. This repeated image-sorting task, involving open-ended discussions on the emotions elicited by each image, allows us to observe a dynamic, iterative collaborative process among the group, where participants achieve a shared goal through incremental steps and consensus. Each group of four participants completed one image-sorting task on the same 28 images and categories. Participants were instructed to work together as a group to reach a consensus on the label of each image, with no time limit for completing the task, allowing participants to engage in deliberate discussion and negotiation. The labels are not placed in a fixed position, allowing participants to organize and use the room space however they want. The task ends once they inform the researcher that they agree with the image groupings.
@@ 3.3.2 Virtual Scene and Cues
As shown in What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality (left, headset’s first-person perspective), all 28 virtual images are scattered around the room, and all six emotion category labels (gray virtual plates) are pasted a little higher than where the images are scattered. Participants can view these images and labels at all times via their headsets. For participants, the cue to start the interaction is not defined by the researcher but decided by each participant which image they want to discuss with the group to sort. This lack of structure in cues is by design, as our goal is to observe open and free collaborations without participants taking turns or being directed by the flow of the virtual scene designed by the researcher. The participants are assumed to take the cue for virtual interaction from the other three group members as shown in What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality (right), where all four participants are physically close, probably examining the same image and discussing the final label.
@@ 3.3.3 Interaction and Feedback
To sort an image, the participants were asked to physically move the virtual image near the virtual label. Once the image is pasted close to the label, the image is recorded as sorted. Participants used a point-and-drag near-interaction motion with the grip buttons on their left or right controllers to move an image to the corresponding label. They pointed at an image, pressed and held the grip button to drag it, and released it to place it. The image followed the controller’s pointer and locked into position when released. Participants could only grab objects within reach and were instructed to release the grip button to secure an image in place. Only one participant could move an image at a time, but multiple images could be held simultaneously. This interaction mimicked real-world object placement and provided immediate feedback once placed.
@@ 3.4 Measurements
This section outlines the measures we used to capture group behavior in the image sorting task, categorized into sensor level, task-related performance metrics, and subjective measures. At the sensor level, we collected data from the headset using custom scripts. We recorded the audio signal from the microphone, positions for location, and eye-tracking data to capture the data for conversation, proximity, and shared attention as described in §2.1.2.
On the task level, we recorded the various virtual object interactions, such as the number of virtual image interactions per participant in a group, to count if a participant grabs a virtual image and then releases it. Throughout the task, we capture the number of label changes per group to count the number of times a particular image changes its label. We also captured distinct groupings for each image per group to count the distinct labels for each image. For instance, if Participant A moved an image to label X, Participant B moved it to label Y, and Participant A moved it back to X, the image would have three label changes and two distinct groupings. We also collected high-level performance metrics, such as completion time. We measured completion time as the time elapsed from when the first image was grabbed to when the last image was placed, indicating the group’s overall time completing the task.
@@
Finally, we collect post-task subjective measures via surveys after the image sorting task, such as presence with the IPQ [51] and PQ [68] questionnaires, cognitive load through NASA-Task Load Index (NASA-TLX) [23]222We are aware of the criticism surrounding the use of NASA-TLX. We applied it to measure “perceived” cognitive workload, rather than actual mental load [35]., and perception of the group behavior through a custom-designed questionnaire. The PQ evaluates factors such as the possibility to act and examine, realism, self-evaluation, and interface quality. The IPQ measures spatial and general presence, realism, and involvement. The presence scores are derived from 33 items (14 IPQ and 19 PQ, the cognitive load score is derived from 5 items from NASA-TLX, and group behavior from 5 items from our custom-designed survey on a 7-point scale. We developed a custom questionnaire to measure participants’ perspectives of how their group interacted, as shown in Table 2. The proposed group behavior characterization questionnaire assesses key aspects such as contribution awareness, shared attention, proximity impact, conversational support, and overall group collaboration. Each question is informed by established research to ensure relevance to our study’s context.
@@ 3.5 Study Procedure
Upon arrival, participants received a detailed information sheet outlining procedures, data collection, and privacy. The researcher also provided a verbal briefing, covering headset interactions, controller gestures, and visual stimuli details such as their color, shape, duration, cues, and feedback mechanism for the image-sorting task. Participants were given ample time to consider their participation in the study and were asked for their verbal consent. Participants then completed a demographic questionnaire covering age, gender, tech familiarity, and familiarity with group members.
@@
Metrics: Mean (), Standard Deviation (), Standard Error (SE), 95% Confidence Interval (CI), 5th/95th Percentiles (P5/P95), Min, and Max.
@@ 4.1 Perceived Group Behavior & Tasks Summary
We summarize the descriptive statistics illustrating participants’ subjective experiences with response variability in Table 3. Metrics such as mean () and standard deviation () PQ () and IPQ () indicate moderate presence levels. We also reported sub-scales: realism (PQ-REAL, IPQ-REAL), possibility to act (ACT), interface quality (IFQUAL), possibility to examine (EXAM), self-evaluation of performance (EVAL), involvement (INV), general presence (GP), and spatial presence (SP). The PQ-REAL subscale (), suggests moderate realism, while the ACT subscale shows a high , indicating strong perceived action capability. INV subscale variability () highlights differing engagement levels. SP () reflects strong spatial awareness. The NASA TLX score () points to a low perceived workload. Group cohesion () and collaboration scored high (), while group proximity varied (), indicating differing closeness perceptions.
@@
| Group | Num of Images Grabbed | Total Num of Image Grabbing | Num of Image Labels Overridden | Total Images Looked At | Completion Time (seconds) | Num of Label Changes |
@@ 4.2 Collaboration via Sensor-based Indicators
This section discusses the participants’ reflections on collaboration, alongside behavioral indicators obtained from sensors and system logs, at both individual and group scales. Findings for shared gaze overlap frequency (indicating visual focus on the same virtual image), proximity duration (measuring time spent in physical proximity during interaction), and speaking proportion and variance (representing turn-taking and conversational balance via total speaking time, frequency, and variance in participation) can be seen in Table 6. Normality checks for subjective collaboration scores were conducted using the Shapiro-Wilk test ( at the group level; - at the individual level), determining the choice of Spearman or Pearson test based on sensor metric normality. Highlighted table rows indicate statistically significant correlations (), offering quantitative insight into how subjective assessments correlate with interaction behavior patterns.
@@
| Sensor-based (Subjective) Metric | Target Normality (Indiv. / Group) | Metric Normality (Indiv. / Group) | Correlation (Indiv. / Group) | p-Value (Indiv. / Group) | Test Used (Indiv. / Group) |
| Shared Gaze Overlap (Shared Attention) | 2.01e-04 / 0.286 | 7.40e-08 / 0.0014 | 0.3646 / 0.4814 | 0.0108 / 0.1131 | Spearman / Spearman |
| Shared Gaze Duration (Shared Attention) | 2.01e-04 / 0.286 | 7.32e-07 / 0.0650 | 0.3027 / 0.2262 | 0.0365 / 0.4796 | Spearman / Pearson |
| Jointly Viewed Images (Shared Attention) | 2.01e-04 / 0.286 | 0.2048 / 0.3740 | -0.0592 / 0.0861 | 0.6893 / 0.7902 | Spearman / Pearson |
| Speaking Proportion (Conversation) | 2.01e-04 / 0.286 | 0.0021 / 0.0097 | 0.0063 / 0.6517 | 0.9659 / 0.0217 | Spearman / Spearman |
| Speaking Variance (Conversation) | 2.01e-04 / 0.286 | 2.71e-05 / 0.0998 | -0.1019 / -0.1117 | 0.4907 / 0.7297 | Spearman / Pearson |
| Proximity Duration (Collaboration) | 2.01e-04 / 0.286 | 7.32e-07 / 0.0650 | 0.3027 / 0.2262 | 0.0365 / 0.4796 | Spearman / Pearson |
| Mean Distance (Collaboration) | 2.01e-04 / 0.286 | 0.5384 / 0.7793 | -0.1913 / -0.0905 | 0.1927 / 0.7797 | Spearman / Pearson |
| Image Grabs (Collaboration) | 2.01e-04 / 0.286 | 3.75e-06 / 0.7734 | -0.1498 / 0.1325 | 0.3094 / 0.6814 | Spearman / Pearson |
| Label Overrides (Collaboration) | 2.01e-04 / 0.286 | 0.0037 / 0.0342 | 0.0881 / 0.1470 | 0.5516 / 0.6485 | Spearman / Spearman |
| Image Grabs Var (Collaboration) | 2.01e-04 / 0.286 | 7.97e-11 / 7.27e-05 | 0.1579 / 0.4921 | 0.2837 / 0.1042 | Spearman / Spearman |
| Label Overrides Var (Collaboration) | 2.01e-04 / 0.286 | 5.52e-04 / 0.3237 | 0.0886 / 0.1727 | 0.5491 / 0.5916 | Spearman / Pearson |

### v3

@@ 2.3 Functional Relational Pathways
We first examine whether sensor-derived behavioral signals correspond to users’ interpretations of collaboration. We call this relationship reflective interaction mapping. While subjective reports are post-hoc and summarized, we analyze objective indicators over task-level windows that match the granularity of users’ reflections. We do not compare specific events to single survey items but assess whether aggregate behavioral patterns (gaze overlap, turn-taking) are reflected in users’ overall experience. This avoids false precision and treats subjective reports as interpretive anchors rather than timestamped outcomes. For example, we ask whether shared gaze events align with a user’s sense of shared attention. Do balanced speech patterns correspond to reported conversational support? This connection motivates our first set of hypotheses, which examine whether subjective perceptions of collaboration are reflected in the observable sensor-based measures.
2.3.2 Trace-to-Performance Alignment (Sensor Perf.)
@@
2.3.3 Subjective Influence on Performance (Subj. Perf.)
Finally, we investigate whether users’ internal experiences correspond to observable group-level performance metrics. We refer to this relationship as a subjective influence on performance. Although not sensor-measured, states such as presence and cognitive load influence how users engage with others and respond to collaborative demands. For example, lower cognitive load may support better coordination [15], leading to more balanced participation and faster task completion. This mapping tests hypotheses that investigate whether subjective experiences (own experience or the group as a whole) align with observable group outcomes.
@@ 2.4 Hypotheses
Together, the three relationships defined by the S2S framework guide the structure of our hypotheses. We will focus our experimental design on testing the following hypotheses:
@@
H3.1: Higher individual perceived presence and perception of contribution is reflected in more frequent shared attention events among group members.
@@ 3.1 Participants
We recruited participants; data from 12 participants were discarded due to technical issues, resulting in participants being included in the analysis. Participants were allowed to form their own groups, or, if preferred, the research team randomly assigned them to a group. In total, 12 groups completed the study. In similar research on collaborative behavior involving small groups, the group is typically defined as having three or more members [65]; in our study, we formed a group with 4 members. Prior studies have revealed patterns of conversation, interaction, and coordination within groups of four in both desktop and virtual reality environments [73]. By adopting a group size of four participants, we aimed to create a richer collaborative environment, capturing the complexities of group behavior not as evident in smaller groups [3, 41]. Participants’ ages ranged from 21 to 42, with a mean age of 24. A summary of the participant demographics can be seen in Table 1. All participants provided verbal informed consent. Each participant had normal or corrected-to-normal vision. The University of California, Irvine Institutional Review Board approved the study.
@@ 3.2 Materials
To capture sensor data on user interactions, we used the Meta Quest Pro [37]. The integrated sensors enabled audio recording, eye-tracking, six-degrees-of-freedom (6 DoF) simultaneous localization and mapping (SLAM) tracking, and MR capabilities. The collaborative app is built in Unity using Meta XR APIs to log audio, gaze, location, and object interactions [36]. For precise manipulation of virtual objects, we use the controller integrated with the Meta Quest Pro [37]. The interaction recording and tracking are validated by others [47, 39]. Participants were invited to a shared lab room with a designated space cleared of materials to minimize distractions. They were informed that they could move freely within this area during the task.
@@ 3.3 Experimental Task
This section summarizes the cue, interaction, and feedback of our collaborative image-sorting task. The study was deliberately designed around a single collaborative task under one shared condition, without varying levels of group stressors or task types. Our goal was not to compare multiple experimental conditions but to examine the richness of group behavior in a naturalistic, unconstrained collaborative setting. Prior work has shown that tightly controlled comparisons can obscure the variability and fluidity of real-time collaboration in MR environments [73]. Instead, we focused on collecting high-resolution behavioral data and subjective reflections during a consistent, shared experience across groups. Each group completed one image-sorting task using the same images and categories. Participants were instructed to work together to reach a consensus on the grouping of each image.
@@ 3.3.1 Primary Task
Participants sorted 28 images from the OASIS dataset [31], which includes 900 validated images rated by 822 participants for pleasantness and arousal. Images were selected to span diverse emotions while avoiding distressing content. Participants sorted the selected images into one of six emotion categories randomly chosen from Russell’s circumplex model of affect [50]: angry, bored, relaxed, tense, pleased, or frustrated. Image sorting tasks have been shown to foster decision-making, communication, and social coordination by building shared mental models and group alignment [10]. We focus on this collaborative task marked by asynchronous, flexible participation, where participants can contribute and modify inputs independently. This repeated image-sorting task, involving open-ended discussions on the emotions evoked by each image, allows us to observe a dynamic, iterative, and collaborative process among the group, where participants achieve a shared goal through incremental steps and consensus. Each group of four participants completed one image-sorting task on the same 28 images and categories. Participants were instructed to work together as a group to reach a consensus on the label of each image, with no time limit for completing the task, allowing participants to engage in deliberate discussion and negotiation. The labels are not placed in a fixed position, allowing participants to organize and utilize the room space as they see fit. The task ends once they inform the researcher that they agree with the image groupings.
@@ 3.3.2 Virtual Scene and Cues
As shown in What Sensors See, What People Feel: An Exploratory Study of Subjective Collaboration Perception in Mixed Reality (left, headset’s first-person perspective), all 28 virtual images are scattered around the room, and all six emotion category labels (gray virtual plates) are pasted a little higher than where the images are scattered. Participants can view these images and labels at all times via their headsets. For participants, the cue to start the interaction is not defined by the researcher but decided by each participant, which image they want to discuss with the group to sort. This lack of structure in cues is by design, as our goal is to observe open and free collaborations without participants taking turns or being directed by the flow of the virtual scene designed by the researcher. The participants are assumed to take the cue for virtual interaction from the other three group members as shown in What Sensors See, What People Feel: An Exploratory Study of Subjective Collaboration Perception in Mixed Reality (right), where all four participants are physically close, probably examining the same image and discussing the final label.
@@ 3.3.3 Interaction and Feedback
To sort an image, the participants were asked to physically move the virtual image near the virtual label. Once the image is pasted close to the label, the image is recorded as sorted. Participants used a point-and-drag near-interaction motion with the grip buttons on their left or right controllers to move an image to the corresponding label. They pointed at an image, pressed and held the grip button to drag it, and released it to place it. The image followed the controller’s pointer and locked into position when released. Participants could only grab objects within reach and were instructed to release the grip button to secure an image in place. Only one participant could move an image at a time, but multiple images could be held simultaneously. This interaction mimicked real-world object placement and provided immediate feedback once placed. Current MR networking protocols restrict virtual object ownership to a single user to maintain state consistency across headsets. This technical constraint naturally facilitated a “negotiate-then-act” workflow, where group consensus preceded individual execution.
@@ 3.4 Measurements
This section outlines the measures we used to capture group behavior in the image sorting task, categorized into sensor-level, task-related performance metrics, and subjective measures. At the sensor level, we collected data from the headset using custom scripts. We recorded the audio signal from the microphone, positions for location, and eye-tracking data to capture the data for conversation, proximity, and shared attention as described in §2.1.2.
At the task level, we recorded various virtual object interactions, such as the number of virtual image interactions per participant in a group, to determine if a participant grabs a virtual image and then releases it. Throughout the task, we capture the number of label changes per group to count the number of times a particular image changes its label. We also captured distinct groupings for each image, per group, to count the distinct labels for each image. For instance, if Participant A moved an image to label X, Participant B moved it to label Y, and Participant A moved it back to X, the image would have three label changes and two distinct groupings. We also collected high-level performance metrics, such as completion time. We measured completion time as the time elapsed from when the first image was grabbed to when the last image was placed, indicating the group’s overall time completing the task.
@@
Finally, we collect post-task subjective measures via surveys after the image sorting task, such as presence with the IPQ [52] and PQ [70] questionnaires, cognitive load through NASA-Task Load Index (NASA-TLX) [23]222Despite criticism, we use NASA-TLX to measure “perceived” cognitive workload, rather than actual mental load [35]., and perception of group behavior through a custom-designed questionnaire. The PQ evaluates factors such as the possibility to act and examine, realism, self-evaluation, and interface quality. The IPQ measures spatial and general presence, realism, and involvement. The presence scores are derived from 33 items (14 IPQ and 19 PQ, the cognitive load score is derived from 5 items from NASA-TLX, and group behavior from 5 items from our custom-designed survey on a 7-point scale. We developed a custom questionnaire to measure participants’ perspectives on how their group interacted, as shown in Table 2. The proposed group behavior characterization questionnaire assesses key aspects such as contribution awareness, shared attention, proximity impact, conversational support, and overall group collaboration. Each question is informed by established research to ensure relevance to our study’s context.
@@ 3.5 Study Procedure
Upon arrival, participants received a detailed information sheet outlining procedures, data collection, and privacy. The researcher also provided a verbal briefing, covering headset interactions, controller gestures, and visual stimulus details, such as their color, shape, duration, cues, and feedback mechanisms, for the image sorting task. Participants were given ample time to consider their participation in the study and were asked for their verbal consent. Participants then completed a demographic questionnaire covering age, gender, tech familiarity, and familiarity with group members.
@@ 4.1 Perceived Group Behavior & Tasks Summary
We summarize the descriptive statistics illustrating participants’ subjective experiences with response variability in Table 3. Metrics such as mean () and standard deviation () PQ () and IPQ () indicate moderate presence levels. We also reported sub-scales: realism (PQ-REAL, IPQ-REAL), possibility to act (ACT), interface quality (IFQUAL), possibility to examine (EXAM), self-evaluation of performance (EVAL), involvement (INV), general and spatial presence (GP, SP). The PQ-REAL subscale () suggests moderate realism, while the ACT shows a high , indicating strong perceived action capability. INV subscale variability () highlights differing engagement levels. SP () reflects strong spatial awareness. The NASA-TLX score () indicates a low perceived workload. Group cohesion () and collaboration () scored high, while group proximity () varied, indicating differing perceptions of closeness.
@@
| Group | Num of Images Grabbed | Total of Image Grabbing | Num of Image Labels Overridden | Total Images Looked At | Completion Time (seconds) | Num of Label Changes |
@@ 4.2 Collaboration via Sensor-based Indicators
This section discusses the participants’ reflections on collaboration, alongside behavioral indicators obtained from sensors and system logs, at both individual and group scales. Findings for shared gaze overlap frequency (indicating visual focus on the same virtual image), proximity duration (measuring time spent in physical proximity during interaction), and speaking proportion and variance (representing turn-taking and conversational balance via total speaking time, frequency, and variance in participation) can be seen in Table 6. Normality checks for subjective collaboration scores were conducted using the Shapiro-Wilk test ( at the group level; - at the individual level), determining the choice of Spearman or Pearson test based on sensor metric normality. For the rest of the paper, to account for multiple hypothesis testing, we applied the Benjamini-Hochberg procedure to calculate False Discovery Rate (FDR) q-values. We report both uncorrected and corrected values. Given the exploratory nature of our study, we treat results where but as promising directional trends for future confirmatory research, while results where are considered statistically robust findings within our S2S framework. Highlighted table rows indicate statistically significant correlations (), offering quantitative insight into how subjective assessments correlate with interaction behavior patterns.
@@
| Sensor-based (Subjective) Metric | Target Normality (Indiv. / Group) | Metric Normality (Indiv. / Group) | Correlation (Indiv. / Group) | p-Value (Indiv. / Group) | q-Value (FDR) (Indiv. / Group) | Test Used (Indiv. / Group) |
| Shared Gaze Overlap (Shared Attention) | 2.01e-04 / 0.286 | 7.40e-08 / 0.0014 | 0.3646 / 0.4814 | 0.0108 / 0.1131 | 0.2376 / 0.4976 | Spearman / Spearman |
| Shared Gaze Duration (Shared Attention) | 2.01e-04 / 0.286 | 3.10e-07 / 0.0059 | 0.0782 / 0.3367 | 0.5972 / 0.2845 | 0.8278 / 0.7563 | Spearman / Spearman |
| Jointly Viewed Images (Shared Attention) | 2.01e-04 / 0.286 | 0.2048 / 0.3740 | -0.0592 / 0.0861 | 0.6893 / 0.7902 | 0.8278 / 0.8278 | Spearman / Pearson |
| Speaking Proportion (Conversation) | 2.01e-04 / 0.286 | 0.0021 / 0.0097 | 0.0063 / 0.6517 | 0.9659 / 0.0217 | 0.9659 / 0.2387 | Spearman / Spearman |
| Speaking Variance (Conversation) | 2.01e-04 / 0.286 | 2.71e-05 / 0.0998 | -0.1019 / -0.1117 | 0.4907 / 0.7297 | 0.8278 / 0.8278 | Spearman / Pearson |
| Proximity Duration (Collaboration) | 2.01e-04 / 0.286 | 7.32e-07 / 0.0650 | 0.3027 / 0.2262 | 0.0365 / 0.4796 | 0.2677 / 0.8278 | Spearman / Pearson |
| Mean Distance (Collaboration) | 2.01e-04 / 0.286 | 0.5384 / 0.7793 | -0.1913 / -0.0905 | 0.1927 / 0.7797 | 0.7066 / 0.8278 | Spearman / Pearson |
| Image Grabs (Collaboration) | 2.01e-04 / 0.286 | 3.75e-06 / 0.7734 | -0.1498 / 0.1325 | 0.3094 / 0.6814 | 0.7563 / 0.8278 | Spearman / Pearson |
| Label Overrides (Collaboration) | 2.01e-04 / 0.286 | 0.0037 / 0.0342 | 0.0881 / 0.1470 | 0.5516 / 0.6485 | 0.8278 / 0.8278 | Spearman / Spearman |
| Image Grabs Var (Collaboration) | 2.01e-04 / 0.286 | 7.97e-11 / 7.27e-05 | 0.1579 / 0.4921 | 0.2837 / 0.1042 | 0.7563 / 0.4976 | Spearman / Spearman |
| Label Overrides Var (Collaboration) | 2.01e-04 / 0.286 | 5.52e-04 / 0.3237 | 0.0886 / 0.1727 | 0.5491 / 0.5916 | 0.8278 / 0.8278 | Spearman / Pearson |